Classify the following as parabolas, ellipses, hyperbolas or circles.
x²+y²-10x+9=0
4y²+9=x
6x²+10y²=60

Answer by Nate(3500)

x²+y²-10x+9=0 is a circle
4y²+9=x is a parabola
6x²+10y²=60 is an ellipse